OpenStudy (anonymous):

evaluate the polynomial for x=-2 -3x to the 2nd power +9x to the 2nd power -9x-2

jhonyy9 (jhonyy9):

\[-3x ^{2} +9x ^{2}-9x-2\]

jhonyy9 (jhonyy9):

do you realy have tought it the same or -3x wann to be on exponent 3 ?

jhonyy9 (jhonyy9):

so for exponent 3 and for x=-2 this is equal -3(-8) +9(4) -9(-2)-2=24+36+18-2=76